OpenStudy (anonymous):

If I a + 3 I =1 and I b - 2 I = 4, what is the one possible value of I a +b I ? a)-6 b)-4 c)0 d)3 e)6

OpenStudy (jiteshmeghwal9):

|a+3|=1 |b-2|=4 a+3=1 b-2=4 Solve the equation for a & b

OpenStudy (jiteshmeghwal9):

& then put the values in the equation |a+b|=a+b=?

OpenStudy (anonymous):

-2?

OpenStudy (jiteshmeghwal9):

Nope